OpenStudy (anonymous):

in a calometry experiment, .50 kg of a metal @ 100 degrees Celsius is added to .50 L of water @ room temperature in an aluminum calometer cup. The cup has a mass of 250 g. a) if the final temperature of the mixture is 25 degrees Celsius, what is then specific heat of the metal? b) What would be the effect if some water splashed out of the cup when the metal was added?

OpenStudy (anonymous):

You'll need the specific heat of water and of aluminum to answer this question, so look those up if you don't have them.

OpenStudy (anonymous):

specific of water is 1 and aluminum 0.21

OpenStudy (anonymous):

*specific heat rather

OpenStudy (anonymous):

To answer this, we use the formula: \[Q = mC\Delta T\] If we have multiple sources for heat transfer: \[Q = (m_1C_1+m_2C_2+...)\Delta T\] Since heat is energy, and energy is conserved (we're assuming heat doesn't actually leave the calorimeter), then the heat leaving the metal equals the heat that enters the water and the aluminum calorimeter. \[m_{metal}C_{metal}\Delta T_{metal} = (m_{water}C_{water} + m_{aluminum}C_{aluminum})\Delta T_{water+calorimeter}\] Now, "delta T" is the change in temperature: \[\Delta T = T_{final}-T_{initial}\] Substitute for that, plug in your initial temperatures, and solve for the final temperature.

OpenStudy (anonymous):

Oops, sorry. Just find delta T, and solve for C metal. :)

OpenStudy (anonymous):

in ΔT water + calorimeter , what will substitute in that calorimeter?

OpenStudy (anonymous):

uhm,,,what's the specific heat for metal? i can't find it ..

OpenStudy (anonymous):

\[\Delta T_{metal} = 100 - 25 = 75\] \[(m_{water}C_{water}+m_{alum}C_{alum})(T_f-T_i) = m_{metal}C_{metal}(75)\] The water and cup were initially at room temperature. T final will be 25 as well (When in equilibrium, everything is the same temperature). Plug those numbers in, solve for C metal.

OpenStudy (anonymous):

okay,, :)

OpenStudy (anonymous):

I didn't do delta T for the water and cup, because "room temperature" definitions vary :)

OpenStudy (anonymous):

if i use 20 degrees celsius?

OpenStudy (anonymous):

.50L is 500 grams? right?

OpenStudy (anonymous):

That's reasonable, I suppose. It's usually mentioned in class at some point. Yeah, .5 L = 500 g for water

OpenStudy (anonymous):

what would be my initial temp.?100 degrees celsius?

OpenStudy (anonymous):

For the water and calorimeter, your initial temperature is room temperature. It tells you that the final temperature is 25 degrees.

OpenStudy (anonymous):

(500 g (1) + 250g (0.21)(25-20) = 500gg Cmetal (75) ...right?

OpenStudy (anonymous):

Looks good to me (you missed a parenthesis on the left, though)

OpenStudy (anonymous):

in 500g?

OpenStudy (anonymous):

Just after the (0.21), before the (25-20)

OpenStudy (anonymous):

What would be the effect if some water splashed out of the cup when the metal was added?

OpenStudy (anonymous):

what will be my unit in my answer? :)

OpenStudy (anonymous):

Well, if water splashes out, then we have less water absorbing the heat from the metal. The final temperature would then be higher than 25, as the remaining water will have to absorb some more energy to reach thermal equilibrium with the metal.

OpenStudy (anonymous):

The units should be that of a specific heat: cal/g K

OpenStudy (anonymous):

ahhhh,,,thank youu :)

OpenStudy (anonymous):

ahhh ...

OpenStudy (anonymous):

after this,,, 2762.5 = 500g Cmetal (75), whats next?

OpenStudy (anonymous):

transpose 500g on the left?

OpenStudy (anonymous):

Oh, I gave the units for your answer for C. What you have on the left hand side has units of calories (energy). Now, you just solve for C: 500*75*C = 2762.5 \[C = \frac{2762.5}{500*75}\]

OpenStudy (anonymous):

ahhhh,,,gets !!